The equation is written incorrectly. I need to rewrite by inserting 1 group of parentheses in the equation to obtain the given anwer. I must follow the order of operations.
6²/-2²+2= -18

Answer by Edwin McCravy(20060)


6²/-2²+2 = -18

Put parentheses around -2²+2

6²/(-2²+2) = -18

36/(-4+2)  = -18

  36/(-2)  = -18

       -18 = -18

Edwin